Contribution Margin and Contribution Margin Ratio

For a recent year, Wicker Company-owned restaurants had the following sales and expenses (in millions):

Sales $30,100
Food and packaging $11,409
Payroll 7,600
Occupancy (rent, depreciation, etc.) 5,791
General, selling, and administrative expenses 4,400
$29,200
Income from operations $900

Assume that the variable costs consist of food and packaging, payroll, and 40% of the general, selling, and administrative expenses.

a. What is Wicker Company's contribution margin? Round to the nearest million. (Give answer in millions of dollars.)
$ million

b. What is Wicker Company's contribution margin ratio? Round to one decimal place.
%

c. How much would income from operations increase if same-store sales increased by $1,800 million for the coming year, with no change in the contribution margin ratio or fixed costs? Round your answer to the closest million.
$ million

Part 2

Currently, the unit selling price of a product is $240, the unit variable cost is $200, and the total fixed costs are $364,000. A proposal is being evaluated to increase the unit selling price to $270.

a. Compute the current break-even sales (units).
units

b. Compute the anticipated break-even sales (units), assuming that the unit selling price is increased to the proposed $270, and all costs remain constant.
units
Answer #1

All the amounts are in million :

a. Total variable cost = Food and packaging + Payroll + 40% of the general, selling and administrative expenses = $11,409 + $7,600 + ( 40% * $4,400 ) = $20,769

Contribution margin = Sales - Total variable cost = $30,100 - $20,769 = $9,331

b. Contribution margin ratio = Contribution margin / Sales = $9,331 / $30,100 = 31%

c. As there is no change in the contribution margin ratio or fixed costs the increase in income from operations will calculated as follows :

Increase in income from operations = Increase in sales * Contribution margin ratio = $1,800 * 31% = $558

Part 2 :

a. Current break even sales (Units) = Total fixed cost / ( selling price - unit variable cost ) = $364,000 / ( $240 - $200 ) = 9,100 units

b. Anticipated break even sales (Units) = Total fixed cost / ( proposed selling price - unit variable cost ) = $364,000 / ( $270 - $200 ) = 5,200 units